Sigma Corporation applies overhead cost to jobs on the basis of direct labor cost. Job V, which was started and completed during the current period, shows charges of $5,000 for direct materials, $8,000 for direct labor, and $6,000 for overhead on its job cost sheet. Job W, which is still in process at year-end, shows charges of $2,500 for direct materials and $4,000 for direct labor. 


Required: 

1a. Should any overhead cost be applied to Job W at year-end? 

1b. How much overhead cost should be applied to Job W? 


2. How will the costs included in Job W's job cost sheet be reported within Sigma Corporation's financial statements at the end of the year?

  • Raw Materials 

  • Work-in-Process 

  • Finished Goods
Answer #1
1a. Yes

1b. Overhead rate = Overhead/Direct labor = 6000/8000 = 75% of dl cost

Overhead cost = 4000*75%

= 3000

2. Work in Process
Answer #2
Req 1-a
YES, the overheads cost must be applied to Job W
Req 1-b
OH rate as % of DLC
Total Overheads 6000
Divide: Direct Labour cost 8000
OH rate as % of DLC 75% of DLC
Overheads cost applied to WIP:
Direct labour cost incurred on Job 4000
Multiply: OH rate as % of DLC 75%
Overheads cost applied to WIP: 3000
Answer is $ 3,000
Req 2.
Work in process